OpenStudy (anonymous):

For f(x)= 5x+2 find f(-3)

OpenStudy (yanasidlinskiy):

does f = -3??

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

-13

OpenStudy (anonymous):

-13 isnt on my answer choice :( its either 1,-1,9 or -9 :(

OpenStudy (yanasidlinskiy):

yup:) it is -13:)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

hhhhh

OpenStudy (anonymous):

put x=-3 in f(x) f(-3)=5*(-3)+2=-15+2=-13

OpenStudy (yanasidlinskiy):

you may have written the question wrong...check it again:)

OpenStudy (anonymous):

oh damn it! your right, yeah i go the wrong question mixed up :)

OpenStudy (anonymous):

okay how about For f(x)= 5x+4 find f(-1)

OpenStudy (yanasidlinskiy):

what is it:)??

OpenStudy (anonymous):

-13 haha

OpenStudy (yanasidlinskiy):

now that would equal -9

OpenStudy (yanasidlinskiy):

just kiddin it would equal 4..my bad

OpenStudy (anonymous):

omg hahahah i was about to put -9 down, thanks

OpenStudy (anonymous):

wwaitt no i dont have 4 down on there lol

OpenStudy (yanasidlinskiy):

ok..then..what is it i don't get your question..check if its right:)maybe thats the problem again.....LOL!!!

OpenStudy (anonymous):

hahaha no i just checked it this time xDD! -9 is an answer choice so ill just put that down

OpenStudy (yanasidlinskiy):

ok..then:D.....i just hope you don't get it wrong:)